### v1

@@ 3 Algorithm
and then obtain the pairwise mirror distance matrix $\mathbf{D}^{\star}\in\mathbb{R}^{m\times m}$. By Theorem 2.6 and Eq. (2), $\mathbf{D}^{\star}$ exhibits a distinct block structure corresponding to the cluster assignments, i.e.,
(3) $\mathbf{D}^{\star}_{i,j}\left\{\begin{aligned} &=0\text{ if }Y_{i}=Y_{j},\\ &\geq c\text{ if }Y_{i}\neq Y_{j}\end{aligned}\right.$
@@
From the observed dynamic adjacency matrices $\{\mathbf{A}^{(i)}_{t}\}_{i\in[m],t\in[T]}$, we can get an estimate of $\mathbf{D}^{\star}$, and this estimate $\hat{\mathbf{D}}^{\star}$ also has the similar block structure. Then the hierarchical clustering algorithm [39, 52] can be applied to the distance matrix $\hat{\mathbf{D}}^{\star}$ to restore the clusters of dynamic networks. To obtain the estimated pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}$ from observed $\{\mathbf{A}^{(i)}_{t}\}_{i\in[m],t\in[T]}$, we first compute the consistent estimates for the underlying unknown latent position matrices $\{\mathbf{X}^{(i)}_{t}\}_{i\in[m],t\in[T]}$ by the rank-$d$ eigendecomposition of the adjacency matrices [48, 47]. More specifically, each $\hat{\mathbf{X}}^{(i)}_{t}$ is the adjacency spectral embedding obtained from its respective adjacency matrix $\mathbf{A}^{(i)}_{t}$.
Definition 3.1 (Adjacency spectral embedding).
@@
0:  $m$ dynamic networks for $T$ time points $\{\mathbf{A}^{(i)}_{t}\in\mathbb{R}^{n\times n}\}_{i\in[m],t\in[T]}$; an embedding dimension $d$ for networks; a dimension $r$ for mirror/low-rank configuration for the distance matrices; a cluster number $K$. Step 1 Constructing the mirror for each dynamic network $i\in[m]$:
  1. 1.
    For each time point $t\in[T]$, compute estimated vertex latent position matrix $\hat{\mathbf{X}}^{(i)}_{t}=\hat{\mathbf{U}}^{(i)}_{t}|\hat{\bm{\Lambda}}^{(i)}_{t}|^{1/2}\in\mathbb{R}^{n\times d}$, where the diagonal matrix $\hat{\bm{\Lambda}}^{(i)}_{t}\in\mathbb{R}^{d\times d}$ contains the $d$ leading eigenvalues of $\mathbf{A}^{(i)}_{t}$ and $\hat{\mathbf{U}}^{(i)}_{t}\in\mathbb{R}^{n\times d}$ contains the corresponding eigenvectors;
  2. 2.
    Construct the estimated distance matrix $\hat{\mathbf{D}}^{(i)}\in\mathbb{R}^{T\times T}$ for the $T$ time points with entries $\hat{\mathbf{D}}^{(i)}_{t_{1},t_{2}}=\frac{1}{\sqrt{n}}\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\hat{\mathbf{X}}^{(i)}_{t_{1}}\mathbf{O}-\hat{\mathbf{X}}^{(i)}_{t_{2}}\|_{F}$ for all $t_{1},t_{2}\in[T]$.
  3. 3.
    Apply CMDS to compute the estimated mirror matrix $\hat{\mathbf{M}}^{(i)}\in\mathbb{R}^{T\times r}$ based on $\hat{\mathbf{D}}^{(i)}$.
Step 2 Clustering dynamic networks according to pairwise mirror distance:
  1. 1.
    Construct the estimated pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}\in\mathbb{R}^{m\times m}$ for the $m$ dynamic networks with $\hat{\mathbf{D}}^{\star}_{i,j}=\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}^{(i)}\mathbf{O}-\hat{\mathbf{M}}^{(j)}\|_{F}$ for all $i,j\in[m]$.
  2. 2.
    Apply hierarchical clustering algorithm to $\hat{\mathbf{D}}^{\star}$ with cluster number $K$ to obtain estimated cluster labels $\{\hat{Y}_{i}\}_{i\in[M]}$.
0:  The estimated cluster labels $\{\hat{Y}_{i}\}_{i\in[M]}$.
@@
In practice, we can choose the dimensions $d$ and $r$ by looking at the eigenvalues of $\{\mathbf{A}^{(i)}_{t}\}$ and $\{\mathbf{D}^{(i)}\}$. A ubiquitous and principled method is to examine the so-called scree plot and look for “elbow” and “knees” defining the cut-off between the top (signal) $d$ or $r$ dimensions and the noise dimensions. [59] provides an automatic dimensionality selection procedure to look for the “elbow” by maximizing a profile likelihood function. [23] suggests another universal approach to rank inference via residual subsampling for estimating the rank. We can also the determine by eigenvalue ratio test [5] or by empirical distribution of eigenvalues [40]. Note that the hierarchical clustering algorithm does not require a predefined number of clusters and can provide a dendrogram, a tree-like chart that illustrates the sequences of merges or splits of clusters. When two clusters are merged, the dendrogram connects them with a line, and the height of the connection represents the distance between those clusters. The optimal number of clusters $K$ can be chosen based on hierarchical structure of the dendrogram, and then be used to obtain clustering result $\{\hat{Y}_{i}\}$.
@@
Note that Algorithm 1 is inherently a distributed algorithm and can benefit from the advantages of distributed algorithms [20, 24, 15]. The network $\{\mathbf{A}^{(i)}_{t}\}$ can be stored on separate local machines. Each local node can perform Step 1 in Algorithm 1 to obtain the $n\times d$ estimated latent position matrix $\hat{\mathbf{X}}^{(i)}_{t}$ and send it to a central node. The central node collects $\{\hat{\mathbf{X}}^{(i)}\}$ and then proceeds with the subsequent steps. With this distributed procedure, the central node does not need the capability to store enormous amounts of data, which is especially important considering that real-world networks may contain a vast number of vertices. And between nodes, only relatively small-scale matrices need to be transferred instead of the raw data, so that the communication cost has been significantly reduced, and the privacy of the raw information has been protected.
@@ 4 Theoretical Results
We now investigate the theoretical properties of our proposed clustering algorithm for $m$ dynamic networks $\{G^{(i)}\}_{i\in[m],t\in[T]}$, each with $n$ common vertices over $T$ time points and cluster structures as described in Section 2. Our main result, presented in Theorem 4.3, establishes the error rate of the estimated pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}$, demonstrating that $\hat{\mathbf{D}}^{\star}$ closely resembles $\mathbf{D}^{\star}$ and thus also captures the block structure corresponding to the clustering assignments, thereby validating the effectiveness of DNCMD. Theorem 4.3 is derived based on the estimation error results for the estimated distance matrix $\hat{\mathbf{D}}^{(i)}$ and the corresponding estimated mirror matrix $\hat{\mathbf{M}}^{(i)}$ in Theorem 4.1 and Theorem 4.2 for each dynamic network $i\in[m]$.
Theorem 4.1.
For a fixed $i\in[m]$, consider a dynamic network $\{G^{(i)}_{t}\}_{t\in[T]}$ as defined in Section 2. Suppose that 1) $n\rho_{n}\gg\log n$, where $\rho_{n}\in(0,1]$ is a sparsity factor depending on $n$ such that for each $i\in[m]$ and each $t\in[T]$, $\|\mathbf{P}^{(i)}_{t}\|=\Theta(n\rho_{n})$; 2) $\{\mathbf{P}^{(i)}_{t}\}$ have bounded condition number, i.e. there exists a finite constant $M>0$ such that $\max_{i\in[m],t\in[T]}\frac{\lambda_{1}(\mathbf{P}^{(i)}_{t})}{\lambda_{d}(\mathbf{P}^{(i)}_{t})}\leq M,$ where $\lambda_{1}(\mathbf{P}^{(i)}_{t})$ (resp. $\lambda_{d}(\mathbf{P}^{(i)}_{t})$) denotes the largest (resp. smallest) eigenvalue of $\mathbf{P}^{(i)}_{t}$. We then have
@@
Condition (1) in the statement of Theorem 4.1 is on the sparsity of the graph. We can interpret $n\rho_{n}$ as the growth rate for the average degrees of the graphs $\{\mathbf{A}^{(i)}_{t}\}$ generated from $\{\mathbf{P}^{(i)}_{t}\}$, and Condition (1) requires that the graphs are not too sparse. Condition (2) pertains to the homogeneity of $\mathbf{P}^{(i)}$ and, consequently, also to the homogeneity of the latent positions. This condition can be relaxed; the bound $M$ does not necessarily have to be constant and can depend on $n$. In this case, the error bound for $\hat{\mathbf{D}}^{(i)}$ will involve $M$. We further establish results for the estimated mirror matrix $\hat{\mathbf{M}}^{(i)}$, based on Theorem 4.1.
Theorem 4.2.
Consider the setting of Theorem 4.1 and further suppose that $\lambda_{r}(\mathbf{B}^{(i)})=\omega(Tn^{-1}(n\rho_{n})^{1/2})$ and $\lambda_{r+1}(\mathbf{B}^{(i)})=O(Tn^{-1}(n\rho_{n})^{1/2})$. We then have
@@
(4) $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}^{(i)}\mathbf{O}-\mathbf{M}^{(i)}\|_{F}\lesssim\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda_{r}^{1/2}(\mathbf{B}^{(i)})}$
@@
The condition about the eigenvalues of the doubly centered distance matrix $\mathbf{B}^{(i)}$ in Theorem 4.2 implies that its largest $r$ eigenvalues and their corresponding eigenvectors contain the primary signal. This further indicates that the $r$-dimensional configuration of these latent position matrices preserves the primary signal.
We then analyze the error of the estimated pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}$ in Theorem 4.3 and the resulting clustering performance based on its structure. Theorem 4.3 shows that the clustering result of DNCMD achieves exact recovery and provides a lower bound for the distinct parameter $c$. Here we use the terminology exact recovery when the partition is recovered correctly with high probability.
Theorem 4.3.
Consider the setting of Theorem 4.2 for all $i\in[m]$. We suppose that there exists a parameter $\lambda$ such that $\lambda_{r}(\mathbf{B}^{(i)})\asymp\lambda$ for all $i\in[m]$. We then have
@@
with high probability. We further suppose that the distinct parameter $c>0$ defined in Eq. (2), which distinguishes the mirrors for different clusters, satisfies $c=\omega(\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda^{1/2}})$. We then have
(5) $\hat{\mathbf{D}}^{\star}_{i,j}\left\{\begin{aligned} &<c/2\text{ if }Y_{i}=Y_{j},\\ &>c/2\text{ if }Y_{i}\neq Y_{j}\end{aligned}\right.$
@@
The error bounds in Theorem 4.2 and Theorem 4.3 depend on the eigenvalues of $\{\mathbf{B}^{(i)}\}$, and the scales of $\{\mathbf{B}^{(i)}\}$ may increase with $T$ as $\{\mathbf{B}^{(i)}\}$ are of size $T\times T$. Actually the relationship $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ can be derived under a general case as described in following Proposition 4.4. Recall that $\rho_{n}$ is the parameter indicating the sparsity of the networks or the magnitude of the latent positions, and here we further use a parameter $\tilde{\rho}$ to describe the magnitude of $\mathbf{D}^{(i)}$ or the change of latent positions across time points.
Proposition 4.4 and the subsequent Proposition 4.5 discuss a general case, providing an intuitive understanding of the conditions under which $\hat{\mathbf{M}}^{(i)}$ converges to the true $\mathbf{M}^{(i)}$, $\hat{\mathbf{D}}^{\star}$ converges to the true $\mathbf{D}^{\star}$, and the requirements for DNCMD to achieve exact recovery.
Proposition 4.4.
Consider the setting of Theorem 4.2. Notice $\|\mathbf{X}_{t}^{(i)}\|_{F}^{2}\asymp n\rho_{n}$, and we further suppose there exists a factor $\tilde{\rho}>0$ to describe the magnitude of the change of $\{\mathbf{X}_{t}^{(i)}\}_{t\in[T]}$ such that $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F}^{2}\asymp n\rho_{n}\tilde{\rho}$ for $t_{1}\neq t_{2}$. We then have $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ and thus
@@
In Proposition 4.4 we suppose $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F}^{2}\asymp n\rho_{n}\tilde{\rho}$ for all pairs $t_{1}\neq t_{2}$ just for ease of exposition. The proof can be adapted to allow $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F}^{2}=0$ for a certain percentage of pairs, thereby handling the more general case.
Proposition 4.4 shows that the estimated mirror $\hat{\mathbf{M}}^{(i)}$ converges to the true mirror $\mathbf{M}^{(i)}$ with high probability when the average degree of graphs $n\rho_{n}$ or the change magnitude $\tilde{\rho}$ is large enough, i.e., $(n\rho_{n})\tilde{\rho}^{2}=\omega(1)$, and the number of time points does not grow too fast, i.e. there exists $\ell>0$ such that $T=O(n^{\ell})$. The discussion in Proposition 4.4 intuitively demonstrates under which conditions our estimator of the mirror matrix is asymptotically convergent.
Proposition 4.5.
Consider the setting of Proposition 4.4 for all $i\in[m]$. We then have
@@
with high probability. We further suppose the distinct parameter $c$ satisfies $c=\omega((n\tilde{\rho})^{-1/2})$. Then Eq. (5) for $\hat{\mathbf{D}}^{\star}$ holds, and hierarchical clustering for $\hat{\mathbf{D}}^{\star}$ achieves exact recovery.
Proposition 4.5 shows the exact recovery of DNCMD requires $c=\omega((n\tilde{\rho})^{-1/2})$. Recall that in this setting, for individual mirror matrices, we obtain $\frac{1}{\sqrt{T}}\|\mathbf{M}^{(i)}\|_{F}\asymp(\rho_{n}\tilde{\rho})^{1/2}$, and for pairs of mirrors from two different clusters with $Y_{i}\neq Y_{j}$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}\geq c$. It follows that, the block structure of $\hat{\mathbf{D}}^{\star}$ in Eq. (5) and the resulting exact recovery of DNCMD require that the ratio of the dissimilarity between clusters to the strength of mirrors satisfies
$\frac{\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}}{\frac{1}{\sqrt{T}}\|\mathbf{M}^{(i)}\|_{F}}\gtrsim\frac{c}{(\rho_{n}\tilde{\rho})^{1/2}}=\omega((n\rho_{n})^{-1/2}\tilde{\rho}^{-1}).$
@@
Remark 4.6.
In the discussion about Proposition 4.5, we do not see that a larger number of time points is beneficial to the error rate. The same phenomenon also appears in Proposition 4.4 for the estimation error for mirrors. It does not mean that we do not need to collect the dynamic networks for more time points. Too inadequate data collection can result in a loss of capture of signals. We consider the example in (a) of Fig. 1. If we just collect the data for time points $1,2,\dots,5$, the difference between dynamic network $1$ and $2$ cannot be captured. Adequate data collection is necessary for detecting discrepancies between dynamic networks, i.e. a large enough $T$ is necessary for accurate $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}$. On the other side, if $T$ is large enough to sufficiently detect the discrepancies between dynamic networks, by the theoretical analysis, we do not need to collect too much data. The example in (b) of Fig. 1 is periodic and the data for one cycle can contain enough signals.
Figure 1: Examples for different dynamic networks

### v2

@@ 3 Algorithm
and then obtain the pairwise mirror distance matrix $\mathbf{D}^{\star}\in\mathbb{R}^{m\times m}$. By Theorem 2 and Eq. (2.2), $\mathbf{D}^{\star}$ exhibits a distinct block structure corresponding to the cluster assignments, i.e.,
$\mathbf{D}^{\star}_{i,j}\left\{\begin{aligned} &=0\text{ if }Y_{i}=Y_{j},\\ &\geq c\text{ if }Y_{i}\neq Y_{j}\end{aligned}\right$. (3.1)
@@
From the observed dynamic adjacency matrices $\{\mathbf{A}^{(i)}_{t}\}_{i\in[m],t\in[T]}$, we can get an estimate of $\mathbf{D}^{\star}$, and this estimate $\hat{\mathbf{D}}^{\star}$ also has the similar block structure. Then the hierarchical clustering algorithm Nielsen and Nielsen (2016); Ward Jr (1963) can be applied to the distance matrix $\hat{\mathbf{D}}^{\star}$ to restore the clusters of dynamic networks. To obtain the estimated pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}$ from observed $\{\mathbf{A}^{(i)}_{t}\}_{i\in[m],t\in[T]}$, we first compute the consistent estimates for the underlying unknown latent position matrices $\{\mathbf{X}^{(i)}_{t}\}_{i\in[m],t\in[T]}$ by the rank-$d$ eigendecomposition of the adjacency matrices Tang and Priebe (2018); Sussman et al. (2012). More specifically, each $\hat{\mathbf{X}}^{(i)}_{t}$ is the adjacency spectral embedding obtained from its respective adjacency matrix $\mathbf{A}^{(i)}_{t}$.
Definition 3 (Adjacency spectral embedding).
@@
0:$m$ dynamic networks for $T$ time points $\{\mathbf{A}^{(i)}_{t}\in\mathbb{R}^{n\times n}\}_{i\in[m],t\in[T]}$; an embedding dimension $d$ for networks; a dimension $r$ for mirror/low-rank configuration for the distance matrices; a cluster number $K$.
Step 1: Constructing the mirror for each dynamic network $i\in[m]$
for each dynamic network $i\in[m]$ do
  for each time point $t\in[T]$ do
   Compute $\hat{\mathbf{X}}^{(i)}_{t}=\hat{\mathbf{U}}^{(i)}_{t}|\hat{\bm{\Lambda}}^{(i)}_{t}|^{1/2}\in\mathbb{R}^{n\times d}$ via eigendecomposition of $\mathbf{A}^{(i)}_{t}$
  end for
  Construct distance matrix $\hat{\mathbf{D}}^{(i)}\in\mathbb{R}^{T\times T}$ with entries $\hat{\mathbf{D}}^{(i)}_{t_{1},t_{2}}=\frac{1}{\sqrt{n}}\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\hat{\mathbf{X}}^{(i)}_{t_{1}}\mathbf{O}-\hat{\mathbf{X}}^{(i)}_{t_{2}}\|_{F}$
  Apply CMDS to compute mirror matrix $\hat{\mathbf{M}}^{(i)}\in\mathbb{R}^{T\times r}$ from $\hat{\mathbf{D}}^{(i)}$
end for
Step 2: Clustering dynamic networks
 Construct pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}\in\mathbb{R}^{m\times m}$ with $\hat{\mathbf{D}}^{\star}_{i,j}=\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}^{(i)}\mathbf{O}-\hat{\mathbf{M}}^{(j)}\|_{F}$
 Apply hierarchical clustering to $\hat{\mathbf{D}}^{\star}$ with $K$ clusters to obtain $\{\hat{Y}_{i}\}_{i\in[m]}$
 The estimated cluster labels $\{\hat{Y}_{i}\}_{i\in[m]}$.
@@
In practice, we can choose the dimensions $d$ and $r$ by looking at the eigenvalues of $\{\mathbf{A}^{(i)}_{t}\}$ and $\{\mathbf{D}^{(i)}\}$. A ubiquitous and principled method is to examine the so-called scree plot and look for “elbow” and “knees” defining the cut-off between the top (signal) $d$ or $r$ dimensions and the noise dimensions. Zhu and Ghodsi (2006) provides an automatic dimensionality selection procedure to look for the “elbow” by maximizing a profile likelihood function. Han et al. (2023) suggests another universal approach to rank inference via residual subsampling for estimating the rank. We can also the determine by eigenvalue ratio test Ahn and Horenstein (2013) or by empirical distribution of eigenvalues Onatski (2010). Note that the hierarchical clustering algorithm does not require a predefined number of clusters and can provide a dendrogram, a tree-like chart that illustrates the sequences of merges or splits of clusters. When two clusters are merged, the dendrogram connects them with a line, and the height of the connection represents the distance between those clusters. The optimal number of clusters $K$ can be chosen based on hierarchical structure of the dendrogram, and then be used to obtain clustering result $\{\hat{Y}_{i}\}$.
@@
Note that Algorithm 1 is inherently a distributed algorithm and can benefit from the advantages of distributed algorithms Fan et al. (2019); Hector and Song (2021); Chen et al. (2022). The network $\{\mathbf{A}^{(i)}_{t}\}$ can be stored on separate local machines. Each local node can perform Step 1 in Algorithm 1 to obtain the $n\times d$ estimated latent position matrix $\hat{\mathbf{X}}^{(i)}_{t}$ and send it to a central node. The central node collects $\{\hat{\mathbf{X}}^{(i)}\}$ and then proceeds with the subsequent steps. With this distributed procedure, the central node does not need the capability to store enormous amounts of data, which is especially important considering that real-world networks may contain a vast number of vertices. And between nodes, only relatively small-scale matrices need to be transferred instead of the raw data, so that the communication cost has been significantly reduced, and the privacy of the raw information has been protected.
@@ 4 Theoretical Results
We now investigate the theoretical properties of our proposed clustering algorithm for $m$ dynamic networks $\{G^{(i)}\}_{i\in[m],t\in[T]}$, each with $n$ common vertices over $T$ time points and cluster structures as described in Section 2. Our main result, presented in Theorem 5, establishes the error rate of the estimated pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}$, demonstrating that $\hat{\mathbf{D}}^{\star}$ closely resembles $\mathbf{D}^{\star}$ and thus also captures the block structure corresponding to the clustering assignments, thereby validating the effectiveness of DNCMD. Theorem 5 is derived based on the estimation error results for the estimated distance matrix $\hat{\mathbf{D}}^{(i)}$ and the corresponding estimated mirror matrix $\hat{\mathbf{M}}^{(i)}$ in Theorem 3 and Theorem 4 for each dynamic network $i\in[m]$.
Theorem 3.
[…]
@@
Condition (1) in the statement of Theorem 3 is on the sparsity of the graph. We can interpret $n\rho_{n}$ as the growth rate for the average degrees of the graphs $\{\mathbf{A}^{(i)}_{t}\}$ generated from $\{\mathbf{P}^{(i)}_{t}\}$, and Condition (1) requires that the graphs are not too sparse. Condition (2) pertains to the homogeneity of $\mathbf{P}^{(i)}$ and, consequently, also to the homogeneity of the latent positions. This condition can be relaxed; the bound $M$ does not necessarily have to be constant and can depend on $n$. In this case, the error bound for $\hat{\mathbf{D}}^{(i)}$ will involve $M$. We further establish results for the estimated mirror matrix $\hat{\mathbf{M}}^{(i)}$, based on Theorem 3.
Theorem 4.
Consider the setting of Theorem 3 and further suppose that $\lambda_{r}(\mathbf{B}^{(i)})=\omega(Tn^{-1}(n\rho_{n})^{1/2})$ and $\lambda_{r+1}(\mathbf{B}^{(i)})=O(Tn^{-1}(n\rho_{n})^{1/2})$. We then have
@@
$\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\hat{\mathbf{M}}^{(i)}\mathbf{O}-\mathbf{M}^{(i)}\|_{F}\lesssim\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda_{r}^{1/2}(\mathbf{B}^{(i)})}$ (4.1)
@@
The condition about the eigenvalues of the doubly centered distance matrix $\mathbf{B}^{(i)}$ in Theorem 4 implies that its largest $r$ eigenvalues and their corresponding eigenvectors contain the primary signal. This further indicates that the $r$-dimensional configuration of these latent position matrices preserves the primary signal.
We then analyze the error of the estimated pairwise mirror distance matrix $\hat{\mathbf{D}}^{\star}$ in Theorem 5 and the resulting clustering performance based on its structure. Theorem 5 shows that the clustering result of DNCMD achieves exact recovery and provides a lower bound for the distinct parameter $c$. Here we use the terminology exact recovery when the partition is recovered correctly with high probability.
Theorem 5.
Consider the setting of Theorem 4 for all $i\in[m]$. We suppose that there exists a parameter $\lambda$ such that $\lambda_{r}(\mathbf{B}^{(i)})\asymp\lambda$ for all $i\in[m]$. We then have
@@
with high probability. We further suppose that the distinct parameter $c>0$ defined in Eq. (2.2), which distinguishes the mirrors for different clusters, satisfies $c=\omega(\frac{T^{1/2}(n\rho_{n})^{1/2}}{n\lambda^{1/2}})$. We then have
$\hat{\mathbf{D}}^{\star}_{i,j}\left\{\begin{aligned} &<c/2\text{ if }Y_{i}=Y_{j},\\ &>c/2\text{ if }Y_{i}\neq Y_{j}\end{aligned}\right$. (4.2)
@@
The error bounds in Theorem 4 and Theorem 5 depend on the eigenvalues of $\{\mathbf{B}^{(i)}\}$, and the scales of $\{\mathbf{B}^{(i)}\}$ may increase with $T$ as $\{\mathbf{B}^{(i)}\}$ are of size $T\times T$. Actually the relationship $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ can be derived under a general case as described in following Proposition 1. Recall that $\rho_{n}$ is the parameter indicating the sparsity of the networks or the magnitude of the latent positions, and here we further use a parameter $\tilde{\rho}$ to describe the magnitude of $\mathbf{D}^{(i)}$ or the change of latent positions across time points.
Proposition 1 and the subsequent Proposition 2 discuss a general case, providing an intuitive understanding of the conditions under which $\hat{\mathbf{M}}^{(i)}$ converges to the true $\mathbf{M}^{(i)}$, $\hat{\mathbf{D}}^{\star}$ converges to the true $\mathbf{D}^{\star}$, and the requirements for DNCMD to achieve exact recovery.
Proposition 1.
Consider the setting of Theorem 4. Notice $\|\mathbf{X}_{t}^{(i)}\|_{F}^{2}\asymp n\rho_{n}$, and we further suppose there exists a factor $\tilde{\rho}>0$ to describe the magnitude of the change of $\{\mathbf{X}_{t}^{(i)}\}_{t\in[T]}$ such that $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F}^{2}\asymp n\rho_{n}\tilde{\rho}$ for $t_{1}\neq t_{2}$. We then have $\lambda_{r}(\mathbf{B}^{(i)})\asymp T\rho_{n}\tilde{\rho}$ and thus
@@
In Proposition 1 we suppose $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F}^{2}\asymp n\rho_{n}\tilde{\rho}$ for all pairs $t_{1}\neq t_{2}$ just for ease of exposition. The proof can be adapted to allow $\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F}^{2}=0$ for a certain percentage of pairs, thereby handling the more general case.
Proposition 1 shows that the estimated mirror $\hat{\mathbf{M}}^{(i)}$ converges to the true mirror $\mathbf{M}^{(i)}$ with high probability when the average degree of graphs $n\rho_{n}$ or the change magnitude $\tilde{\rho}$ is large enough, i.e., $(n\rho_{n})\tilde{\rho}^{2}=\omega(1)$, and the number of time points does not grow too fast, i.e. there exists $\ell>0$ such that $T=O(n^{\ell})$. The discussion in Proposition 1 intuitively demonstrates under which conditions our estimator of the mirror matrix is asymptotically convergent.
Proposition 2.
Consider the setting of Proposition 1 for all $i\in[m]$. We then have
@@
with high probability. We further suppose the distinct parameter $c$ satisfies $c=\omega((n\tilde{\rho})^{-1/2})$. Then Eq. (4.2) for $\hat{\mathbf{D}}^{\star}$ holds, and hierarchical clustering for $\hat{\mathbf{D}}^{\star}$ achieves exact recovery.
Proposition 2 shows the exact recovery of DNCMD requires $c=\omega((n\tilde{\rho})^{-1/2})$. Recall that in this setting, for individual mirror matrices, we obtain $\frac{1}{\sqrt{T}}\|\mathbf{M}^{(i)}\|_{F}\asymp(\rho_{n}\tilde{\rho})^{1/2}$, and for pairs of mirrors from two different clusters with $Y_{i}\neq Y_{j}$, we have $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}\geq c$. It follows that, the block structure of $\hat{\mathbf{D}}^{\star}$ in Eq. (4.2) and the resulting exact recovery of DNCMD require that the ratio of the dissimilarity between clusters to the strength of mirrors satisfies
$\frac{\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}}{\frac{1}{\sqrt{T}}\|\mathbf{M}^{(i)}\|_{F}}\gtrsim\frac{c}{(\rho_{n}\tilde{\rho})^{1/2}}=\omega\Big(\frac{(n\tilde{\rho})^{-1/2}}{(\rho_{n}\tilde{\rho})^{1/2}}\Big)=\omega((n\rho_{n})^{-1/2}\tilde{\rho}^{-1}).$
@@
Remark 3.
In the discussion about Proposition 2, we do not see that a larger number of time points is beneficial to the error rate. The same phenomenon also appears in Proposition 1 for the estimation error for mirrors. It does not mean that we do not need to collect the dynamic networks for more time points. Too inadequate data collection can result in a loss of capture of signals. We consider the example in (a) of Figure 2. If we just collect the data for time points $1,2,\dots,5$, the difference between dynamic network $1$ and $2$ cannot be captured. Adequate data collection is necessary for detecting discrepancies between dynamic networks, i.e. a large enough $T$ is necessary for accurate $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}$. On the other side, if $T$ is large enough to sufficiently detect the discrepancies between dynamic networks, by the theoretical analysis, we do not need to collect too much data. The example in (b) of Figure 2 is periodic and the data for one cycle can contain enough signals.
Figure 2: Examples for dynamic networks.